### v1

# dense matter, dark matter, compact star

Abstract
The fact that strange sea quarks are abundant in the nucleons, but without strangeness, is of great importance for understanding the nature of dense matter condensed by the strong interaction, particularly in the context of the “gigantic nucleus” formed by the gravitational collapse of an evolved massive star. We hypothesize that the basic unit of bulk strong matter with the light-flavor symmetry of valence quarks is “strangeon”, which is the counterpart of the nucleon found in atomic nuclei. In addition to strangeon stars (SnSs) with large baryon number of $A\approx 10^{57}$, strange nuggets (SnNs) with $A\gtrsim 10^{10}$ could also exist in the Universe. Both the SnS and the SnN are explained, with attention to their observational evidence.
keywords:
dense matter, dark matter, compact star
[…]
1 Introduction
The motivation of Physics is to understand the material world, phenomenologically, the condensation of basic units. The normal material of atoms/ion-electrons condensed by the electromagnetic (or simply electric) force is well known as the only form of matter until the discovery of Ernest Rutherford in 1911 (Xu, 2024), revealing a novel matter of nucleons condensed by the fundamental strong interaction in the current discourse. For the sake of simplicity, one might refer to the former as electric matter, and the latter as strong matter.
Can strong matter be manifested only in the form of atomic nuclei? The answer was supposed to be “yes”, even pondering the remnants of evolved massive stars, i.e., the “gigantic nucleus” initiated by Lev Landau in 1932, the so-called “neutron sphere”. It is noteworthy that the key point of Landau’s idea of a “neutron star” is to neutralize the constituent elements — creating an almost charge-neutral state for nucleons so that a relatively small amount of electrons  [^1] As an exemplar of fermions following the Fermi-Dirac statistics, an electron must populate a higher energy level if the lower levels are occupied, as is in the case with atoms that have many electrons. With the Heisenberg relation, $\Delta x\cdot\Delta p\sim\hbar$, relating the separation between electrons $\Delta x$ to the electron’s number density $n\sim\Delta x^{-3}$, one can easily have the typical momentum of an electron $\simeq\hbar n^{1/3}$. It is then evident that a small charge-to-mass ratio, $R_{\textbf{em}}$, for nucleons favours a lower and stable state of electron gas. We would therefore think that a gigantic nucleus should be a neutron star because $R_{\textbf{em}}$(neutron) = 0, whereas $R_{\textbf{em}}$(proton) = 1. In this sense, Landau’s neutron star is, in fact, a neutral star (Xu, 2023). Besides the electron gas, nonetheless, there are also nucleons in the system, and the so-called symmetry energy emerges dramatically for a neutron-rich gigantic nucleus. But that’s for later, “Press down the gourd, but the ladle pops up” in a Chinese proverb, Landau (1932) was unaware of the symmetry energy, which was first addressed by Weizsäcker in 1935 Weizsäcker (1935). would contribute negligible kinematic energy.
However, is there an alternative, and possibly better, way to neutralize today (more than 90-years later, especially after the establishment of the standard model of particle physics)? Yes, the strangeness matters! In comparison with the low-energy system of atom (nucleus and electrons), there are abundant quark-pairs inside the high-energy system of nucleon (e.g., the complex proton [^2] This is also simply due to the Heisenberg relation. An atom is a state bound by the electromagnetic force, while a proton by the strong force. This results in a relatively large atom but a smaller proton. According to the Heisenberg relation, the typical kinematic energy of an electron in a hydrogen atom is $\sim 10$ eV, whereas that of a quark in a proton is $E_{\rm scale}\sim 300$ MeV Xu (2018). The fact that the bare masses of all the light-flavour quarks are $<E_{\rm scale}$ (but the bare masses of all the heavy-flavour quarks are $\gg E_{\rm scale}$) means that rich strange-quark pairs do exist in a nucleon even without strangeness. In fact, like any bound-states (e.g., hadrons) else, the proton is an object of unspeakable complexity: filled with the dance of short-lived $uds$-quarks (i.e., the sea quarks) as well as gluons! Nevertheless, the time-averaged number of net $u$-quarks is 2 (i.e., the number of $u$-quarks minus that of $\bar{u}$-quarks), and the number of net $d$-quarks is 1, but the number of net $s$-quarks is 0. We can therefore conclude that the valence quarks in a proton are {$uud$} (or {$u^{2}d^{1}s^{0}$}). ). It is then convenient to consider the charge-neutrality using the triangle illustrated in Figure 1. Given that the charges of $d$- and $s$-quarks are equal, the gray scale of the image in Figure 1 can be used to represent the charge-to-mass ratio, $R_{\textbf{em}}$. Landau (1932) anticipated superficially the transformation of atomic nuclei from point “A” to a gigantic one at point “n”, but with 2-flavoured asymmetry of valence quarks. However, the most beautiful place in the triangle should be in point “s”, with both the charge neutrality and the 3-flavoured symmetry. We refer to the matter at point “s” as “strange matter”, which contains almost equal numbers of $u$, $d$ and $s$ valence quarks.
[…]
Figure 1: Triangle of light-quark flavors ($u$, $d$, and $s$(Xu et al., 2021). A point inside the triangle defines a state with a specific ratio of up, down and strange quarks, as measured by the point’s distance from each of the triangle’s edges. The charge-to-mass ratio of the baryonic component, $R_{\textbf{em}}$, is illustrated by the grey scale. Atomic nuclei are at point “A”, pure neutron matter is at point “n”, pure proton matter is at point “p”, strange quark matter and strangeon matter (SnM) are at point “s”.
Are quarks truly the fundamental building blocks of strange matter? Yes, strange matter is favoured by both the charge neutrality and the flavour symmetry of quarks, but the puzzling non-perturbative nature of the quantum chromo-mechanism (i.e., NQCD) does work in this case. There would then be two significant consequences of NQCD: quark-clustering (grouped in position space) and core-repulsing (hard core). Both these two are phenomenologically understood for atomic nuclei, but they can hardly be explained by NQCD even today. Nevertheless, as an analogy of nucleons in point “A”, one could also expect a nucleon-like unit in point “s” Xu (2003). We name this unit “strangeon”, and are explaining strangeon matter (SnM) in the following sections.
2 An evolving trend during gravitational contraction
Let us begin with a gravitational system (mass $M$ and radius $R$) that has spherical symmetry. It seems that the system would be most stable when $R\rightarrow 0$ (i.e., becoming a black hole), since its gravitational energy is the minimum, $E_{\mathrm{g}}\simeq-GM^{2}/R\rightarrow-\infty$, in this case. However, there is an internal energy of matter, and the system could temporarily reach a stable state during gravitational contraction if the internal energy is comparable to $-E_{\mathrm{g}}$. This is why, besides black holes, both main-sequence and compact stars do exist in the Universe.
The total energy of the system reads
$E(M)=\epsilon\frac{M}{m_{\mathrm{u}}}+E_{\mathrm{g}}\simeq\epsilon\frac{M}{m_{\mathrm{u}}}-\frac{GM^{2}}{R}=(\epsilon-\frac{GM{m_{\mathrm{u}}}}{R})\frac{M}{m_{\mathrm{u}}},$ (1)
where $\epsilon$ is the internal energy averaged per baryon of matter, $m_{\mathrm{u}}c^{2}=931.5$ MeV $\sim 1$ GeV is the atomic mass unit, $M/m_{\mathrm{u}}$ represents the total baryon number, $A$. When the system reaches an equilibrium, its internal energy is comparable to the absolute value of gravitational energy, one then has
$R\approx\frac{GMm_{\mathrm{u}}}{\epsilon}=\frac{GM_{\odot}}{c^{2}}\frac{M}{M_{\odot}}\frac{m_{\mathrm{u}}c^{2}}{\epsilon}\simeq 1.5\,\mathrm{km}\frac{M}{M_{\odot}}\frac{1~{}{\rm GeV}}{\epsilon}.$ (2)
In other words,
$\epsilon\simeq 1~{}{\rm GeV}\frac{M}{M_{\odot}}\frac{1.5\,\mathrm{km}}{R}.$ (3)
As a system with negative heat capacity, the kinematic internal energy, $\epsilon$, increases as the gravity-bound object releases more energy (i.e. as the radius decreases). This fact results in the emergence of various objects, from main-sequence stars (ϵsimilar-toitalic-$\epsilon\sim$ keV) to white dwarfs (ϵsimilar-toitalic-$\epsilon\sim$ MeV) and neutron stars ($\epsilon\sim 10^{2}$ MeV), and eventually black holes ($\epsilon\rightarrow-\infty$), as illustrated in Figure 2. Certainly, we assume that the initial mass of the system is high enough to go toward a black hole, otherwise it would stop at a step of a white dwarf or a neutron star.
[…]
Figure 2: The three equilibrium steps before an object with a high initial mass collapses into a black hole. The three steps are: first, main-sequence stars, which are supported by thermal gas pressure and nuclear power; second, white dwarfs, which are supported by fermionic degeneracy pressure supplied by the electron gas; and third, neutron stars, which are supported by the strong interaction (e.g. hard core) between baryons in the non-perturbative regime. These are ordered by radius from large to small and by internal energy per particle from low to high.
In Figure 2, the first step represents the main-sequence stars. The nucleosynthesis process is ongoing in the core of the main-sequence stars, maintaining a temperature gradient throughout the stars that allows the thermal pressure of the ionized gas to be strong enough to compete with gravity. The thermal kinetic energy of a particle in the core of the stars is $10^{-3}$ MeV (corresponds to the nucleosynthesis temperature, $10^{7}$ K), the radius of a main-sequence star is then $R\simeq 10^{6}\,{\rm km}\cdot M/M_{\odot}$ from Equation 2, for the Sun as an example.
The second step represents the white dwarfs. When low-mass stars (about $0.5\,M_{\odot}\sim 8\,M_{\odot}$) run out of their nuclear energy, their cores will undergo gravitational collapse and become white dwarfs. When the electron gas reaches high density to become a relativistic degenerate Fermi gas, the equilibrium between the star’s self-gravity and the degeneracy pressure will become unstable and therefore there exists a maximum mass $\sim 1.44\,M_{\odot}$, for white dwarfs, called the Chandrasekhar limit (Chandrasekhar, 1931). So, the internal energy in a white dwarf is comparable to the rest energy of the electrons, 0.5 MeV. The radius of a white dwarf is then $R\simeq 10^{3}\,{\rm km}$ from Equation 2.
The third step represents the “neutron stars” [^3] Although these objects are commonly referred to as neutron stars in the literature, they do not necessarily consist of neutrons., which are manifested in the form of pulsar-like compact stars. A typical observed pulsar has $\sim 1.4\,M_{\odot}$ mass and $\sim 10\,$km radius, so the internal energy of a particle is $\epsilon\simeq 10^{2}\,$eV according to Equation 3. Within this energy range, quantum chromodynamics (QCD) is non-perturbative, the strong interaction can not be negligible but hard to calculate (Li et al., 2019). The basic unit of neutron stars may not be deconfined quarks but be hadron-like particles. In addition, the energy scale is enough to excite strange quarks as valence ones, and the matter in the neutron star may have non-zero strangeness. As explained in Fig. 1, the colour-confined, hadron-like strange particle is a strangeon. Therefore, “neutron stars” may actually be strangeon stars (SnSs).
When the energy scale reaches $\gtrsim$ 1 GeV, a (1 $M_{\odot}$)-star would contract to its Schwarzschild radius from Equ. 3. This hints that itinerate quarks could hardly appear in “neutron stars” because the deconfined quarks may only emerge at the energy scale $\epsilon\gtrsim 1\,$GeV, except for massive objects with masses approaching the maximum mass (e.g., hybrid SnS Zhang et al. (2023)). When all other forms of energy cannot compete with gravitational energy, the internal energy of a particle tends towards infinity and the star inevitably collapses to form a black hole.
3 Strangeon stars
From the analogy that non-perturbative strong interaction makes quarks form nucleons, we conjecture that the base unit of cold dense matter inside the pulsar is not quarks, but strangeons. This model was proposed in 2003 (Xu, 2003), and is constantly evolving, e.g. (Lai and Xu, 2009, 2016; Lu et al., 2023; Yuan et al., 2025).
There are a lot of mass-radius observations to support the SnS model. The photospheric radius expansion bursts of 4U 1746-37 constrain the mass-radius range of low-mass pulsars, and strangeon-star equation of state (Lai and Xu, 2009) meets the constraint better than neutron-star equations of state (Li et al., 2015), see 3. The discovery of massive neutron stars with mass $M\gtrsim 2M_{\odot}$ such as PSR J1614-2230 (Demorest et al., 2010), PSR J0740+6620 (Cromartie et al., 2020), and PSR J0952-0607 (Romani et al., 2022) supports a stiffer equation of state. As seen in Figure 3, SnM is stiffer and can reach a $3M_{\odot}$ maximum mass.
[…]
Figure 3: The mass-radius curve calculated using Tolman-Oppenheimer-Volkoff equation and equations of state, This figure is from (Lai et al., 2025). The first five types in the legend is the different kinds of equations of state of SnSs. The MIT bag model is a quark state, and the AP4 is a neutron-proton-electron-muon state.
Furthermore, the SnM state can help us to understand subpulses in some single pulses. Because the base unit of SnM is non-relativistic strangeons, the strong interaction between strangeons is strong enough to make the matter solid. This solid SnM is different from a neutron star’s solid crust because the latter is mainly combined by electromagnetic interaction. So, an SnS could support a higher and longer lasting “zit” (little hill) than a neutron star’s crust could support. These zits may be the spark origin of some highly linear-polarized subpulses (Vivekanand and Radhakrishnan, 1981; Cao et al., 2024; Wang et al., 2024), see Figure 4.
[…]
Figure 4: Schematic diagram of the highly linear-polarized spark signal at certain pulse phases caused by zits in the polar cap of the pulsar. The point µis the magnetic pole, the black circle represents the boundary of the polar cap. Some zits on the line of sight spark occasionally and generate a coherent linear-polarized subpulse. This figure refers to (Cao et al., 2024; Wang et al., 2024).
4 Strangeon nuggets
The SnM can also form small nuggets, called strangeon nuggets (SnNs). The baryon number of smallest stable SnN, i.e. the critical baryon number $A_{\mathrm{c}}$, is uncertain due to the difficulty of computation of the non-perturbative strong interaction. If one considers the role of the electron in breaking the symmetry of $e^{\pm}$ (Xu, 2020), the critical baryon number can be estimated from the Compton wavelength of the electron, $A_{\mathrm{c}}\approx(\lambda_{\mathrm{C}}/r_{\rm{nucleon}})^{3}\approx 10^{10}$. We have only ever observed strong interacting matter with baryon number $A<300$ (atomic nuclei) or $A\approx 10^{57}$ (pulsars). If this assumption holds, the mass spectrum of strong interacting matter can be broadened considerably (Wu et al., 2024; Xu, 2024), see Figure 5.
[…]
Figure 5: The mass spectrum of the black hole, the strong matter, and the electric matter (i.e. the ordinary matter). The minimum mass of the black hole is the Planck mass. The minimum mass of the strong matter depends on the critical baryon number $A_{\mathrm{c}}$, which is uncertain because of the unknown behavior of the non-perturbative strong interaction. This picture is from (Xia et al., 2024).
SnNs with $10^{25}<A<10^{35}$, as quark nuggets, are among the candidates for dark matter. Since the light-three-flavor symmetry in SnM is only slightly broken by the larger strange-quark mass, the charge-to-mass ratio of SnM is low, the electromagnetic signals from SnN are too weak to be observed. SnNs move virially in the universe at a typical speed of about $10^{-3}c$, when they hit Earth, they may not generate a high-energy shower signal, but rather an acoustic shock wave. So, we can detect these non-relativistic nuggets using the acoustic method (Qi and Xu, 2025). Due to the extremely high density of SnNs, their penetration depth in rock or water is very large, even exceeding the radius of the Earth, the calculation results is shown in Figure 6, in which the dimensionless $k$ factor represents the ratio of the assumed interaction cross-section between the SnN and the ordinary matter and the geometric cross-section of the SnN.
[…]
Figure 6: The penetration depth $x$ of the strangeon nugget (SnN) in Earth. The black line, $10^{8.8}\,$cm, is the radius of the Earth. The dimensionless $k$ factor is the ratio of the assumed interaction cross-section between the SnN and the ordinary matter and the geometric cross-section of the SnN. This picture is from (Qi and Xu, 2025).
The initial shock wave is strong but will decay during propagation in the medium due to diffusion, dispersion, dissipation and scattering. The decay can be described as an exponential function. At a critical distance $R_{\rm crit}$, the acoustic signal decay to the level of surrounding noise, $R_{\rm crit}$ is the maximum detectable distance. This distance depends on the state, the temperature and the uniformity of the medium. An estimated result is in Figure 7, $R_{\rm crit}\simeq 30\,$km. More accurate results require on-site acoustic experiments.
[…]
Figure 7: The estimated critical detectable distance $R_{\rm crit}$ of the acoustic SnN signal. This picture is from (Qi and Xu, 2025).
5 Conclusions
We proposed that pulsar-like compact stars could be SnSs if Nature favors the three-flavor symmetry of light quarks during core-clollapse supernovae. With FAST’s polarization observation, non-symmetric sparking is found, which could imply “zits” (i.e., small hills) on the surface of solid SnS. Additionally, SnNs with a baryon number of $A\gtrsim 10^{10}$ could also exist, and we could potentially detect them using an acoustic array in the future. These nuggets, which could manifest in the form of dark matter, may have been produced during the separation of the QCD phases in the early Universe, setting the initial conditions for big bang nucleosynthesis (BBN).
[…]
Conceptualization, Renxin Xu; investigation, Haoyang Qi; writing—original draft preparation, Haoyang Qi; writing—review and editing, Renxin Xu; All authors have read and agreed to the published version of the manuscript.
[…]
This research received no funding.
[…]
The authors declare no conflicts of interest.
[…]
Abbreviations The following abbreviations are used in this manuscript:
SnM Strangeon matter
SnN Strangeon nugget
SnS Strangeon star
[…]